# Ringleader ASGD: The First Asynchronous SGD with Optimal Time Complexity under Data Heterogeneity

### v2

@@ 1 Introduction
A more promising approach is to shift from hardware solutions to algorithmic ones by adopting asynchronous optimization methods. These methods remove the need for synchronization, allowing fast workers to contribute updates without waiting for slower ones (Tsitsiklis et al., 1986; Recht et al., 2011; Agarwal & Duchi, 2011; Dean et al., 2012; Li et al., 2014). Despite their appeal, asynchronous methods are more difficult to analyze. In particular, a meaningful analysis would require studying time to convergence, rather than iteration complexity only. While iteration complexity is the traditional metric in optimization, it does not necessarily reflect real-world training speed in parallel settings: a method that performs more iterations may finish faster in wall-clock time if those iterations can be computed without waiting for slow workers. This distinction raises a fundamental question: among all parallel methods, which ones are provably fastest in theory? To make this question precise, we restrict our attention to smooth nonconvex problems and to stochastic first-order methods, encompassing algorithms with or without synchronization. This will be the only setting considered in this paper.
@@
Each column indicates whether a method satisfies the following desirable properties: Optimal: achieves the theoretical lower bound derived by Tyurin & Richtárik (2024) for parallel first-order stochastic methods in heterogeneous data setting. No sync.: does not require synchronization and is therefore asynchronous. No idle workers: all workers remain busy without waiting, so computational resources are fully utilized. No discarded work: no computation is wasted, and no worker is stopped mid-computation. Our new method, Ringleader ASGD, is the first asynchronous method to achieve optimal time complexity, while also ensuring full resource utilization (no idle workers) and no discarded computations / work.
@@
Naive Minibatch SGD
(Section 3.1)
$\frac{L_{f}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{n}\frac{\sigma^{2}}{n\varepsilon}\right)$
IA2SGD
(Wang et al., 2025)
(Appendix D)
$\frac{L_{\max}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{n}\frac{\sigma^{2}}{n\varepsilon}\right)$ (†)
Malenia SGD
(Tyurin & Richtárik, 2024)
$\frac{L_{f}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{\mathrm{avg}}\frac{\sigma^{2}}{n\varepsilon}\right)$
Ringleader ASGD (new)
(Algorithm 1; Section 5.2)
$\frac{L\Delta}{\varepsilon}\left(\tau_{n}+\tau_{\mathrm{avg}}\frac{\sigma^{2}}{n\varepsilon}\right)$ (‡)
@@
Recently, Tyurin & Richtárik (2024) studied this very regime, where they derived lower bounds. They then proposed two algorithms: Rennala SGD, designed for the homogeneous data setting, where all workers draw samples from the same distribution, and Malenia SGD, for the heterogeneous data setting, where data distributions differ across workers. They showed that both methods are optimal—achieving the lower bounds—and, perhaps surprisingly, both are synchronous (they periodically synchronize the workers). The key idea in both is to fully utilize the available computational resources by keeping workers continuously busy: each worker computes independently, and synchronization occurs only after a sufficient number of gradient computations have been accumulated.
At first, the result of Tyurin & Richtárik (2024) suggested a rather pessimistic outlook for asynchronous methods: despite their practical appeal, they showed that existing asynchronous methods are not optimal and that the method achieving the lower bound is synchronous. This created the view that optimality is inherently tied to synchronization. However, this view was overturned by Maranjyan et al. (2025d), who, in the homogeneous data setting, introduced Ringmaster ASGD—the first asynchronous SGD method to achieve the same optimal time complexity as the synchronous Rennala SGD. Although both methods share the same theoretical guarantees, Ringmaster ASGD can be faster than Rennala SGD in practice, since it avoids synchronization and benefits from more frequent updates.
@@
To the best of our knowledge, the only method proven to be optimal in the heterogeneous data setting is the synchronous algorithm Malenia SGD of Tyurin & Richtárik (2024), which, notably, does not rely on similarity assumptions. However, synchronization is a major bottleneck in practice: although synchronous and asynchronous methods can share the same theoretical complexity, asynchronous methods are often faster in practice because they avoid costly synchronization and benefit from more frequent updates, as demonstrated in the homogeneous case by Maranjyan et al. (2025d).
This raises a fundamental question: Is it possible to design an asynchronous method that requires no similarity assumptions while still achieving optimal time complexity? In this paper, we answer this question affirmatively by introducing Ringleader ASGD, the first asynchronous SGD method that achieves optimal time complexity [^1] Throughout the paper, we refer to our method as optimal. Formally, this holds whenever the constant LL—associated with our new smoothness-type assumption (Section 2.3)—is at most a constant factor larger than the smoothness constant $L_{f}$ used in the derived lower bounds (Tyurin & Richtárik, 2024). See Table 1 for details. in the heterogeneous data setting. Importantly, Ringleader ASGD attains this without relying on restrictive similarity assumptions.
@@ 1.1 Contributions
  • Optimal asynchronous SGD under data heterogeneity. We prove that Ringleader ASGD (Algorithm 1) is, to the best of our knowledge, the first asynchronous method in the heterogeneous setting under the fixed computation model (2) that matches the lower bounds for parallel methods established by Tyurin & Richtárik (2024) (Table 1). Crucially, it does not rely on any similarity assumptions across clients’ datasets.
@@
  • Parameter-free design. In contrast to the optimal synchronous method Malenia SGD (Tyurin & Richtárik, 2024), which requires prior knowledge of the gradient variance bound and target accuracy, our method operates in the fixed computation time model without such parameters (except for the stepsize, needed only to match the optimal theoretical rate). This makes it far more practical for real-world deployments, where these quantities are typically unknown or difficult to estimate. The same parameter-free improvement can also be extended to Malenia SGD, as we discuss in Appendix E.
  • Universal computation model. In Appendix B, we extend our analysis beyond the fixed computation time model to the general setting of arbitrarily varying computation times, accommodating virtually any computational behavior, including stochastic or adversarial patterns, while retaining optimality time complexity.
@@ 2.1 Worker Heterogeneity Model
We first focus on the case where workers have constant computation speeds, as this setting is more intuitive and serves as a foundational model for understanding the dynamics of asynchronous distributed optimization. The extension to arbitrary computation times is presented in Appendix B.
@@
We assume communication is infinitely fast (taking 0 seconds), both from workers to the server and from the server to workers [^3] Alternatively, one could define $\tau_{i}$ as the time required for a worker to both compute a gradient and communicate it to the server, while keeping server-to-worker communication infinitely fast. Our upper bounds would still hold under this formulation, but the lower bounds would no longer apply, so we use the simpler model.. This is a modeling choice—arguably the simplest one—and has been the standard assumption in prior work (Mishchenko et al., 2022; Koloskova et al., 2022; Tyurin & Richtárik, 2024; Maranjyan et al., 2025d), even if not always stated explicitly. A related study by Tyurin et al. (2024b) considers the case where communication is non-negligible and proposes techniques to address it.
@@ 2.3 Assumptions
By convention, $L_{\phi}$ denotes the smallest such constant. Note that Section 2.3 is stronger than requiring $f$ itself to be $L_{f}$–smooth, yet weaker than all $f_{i}$ being $L_{f_{i}}$–smooth. The following lemma establishes the relation among the constants $_{f}, LL$, and $L_{f_{i}}$. {boxedlemma}[Smoothness Bounds; Proof in Section C.1] Suppose Section 2.3 holds with constant $L>0$. Then $f$ is $L_{f}$–smooth with $L_{f}\leq L$. Moreover, if each $f_{i}$ is $L_{f_{i}}$–smooth, then Section 2.3 is satisfied, and we have
@@ 3 Background and Motivation
In this section, we review relevant existing methods in distributed optimization and discuss their limitations to motivate our algorithmic design. We begin with Naive Minibatch SGD as the canonical synchronous baseline, then consider Malenia SGD (Tyurin & Richtárik, 2024), the first synchronous SGD method to attain optimal time complexity. We then turn to the challenges of asynchronous approaches, focusing on the recent IA2SGD (Wang et al., 2025) and its limitations. Finally, we outline how these limitations can be addressed and introduce the core idea behind our algorithm.
@@ 3.2 Malenia SGD
Malenia SGD (Tyurin & Richtárik, 2024) addresses the straggler problem of Naive Minibatch SGD by ensuring continuous worker utilization, rather than waiting for the slowest worker in each iteration. However, it is fundamentally a Minibatch SGD algorithm: in every iteration, it constructs an unbiased gradient estimator from multiple gradients and then performs a synchronous update. The key distinction lies in how the batch is collected—Malenia SGD gathers gradients asynchronously, allowing potentially multiple contributions from the same worker within a single iteration, unlike Naive Minibatch SGD.
@@
as shown by Tyurin & Richtárik (2024). The key improvement over Naive Minibatch SGD is that the variance term $\sigma^{2}$ is now multiplied by $\tau_{\text{avg}}$ instead of $\tau_{n}$. Since $\tau_{\text{avg}}\ll\tau_{n}$ in computationally heterogeneous environments, Malenia SGD can potentially provide substantial speedup in highly heterogeneous regimes.
@@ 3.4 IA2SGD
The iteration complexity of this algorithm was established by Wang et al. (2025), and by a straightforward conversion to the fixed computation time model (see Appendix D), we obtain the corresponding time complexity
@@ 4.4 Comparison to Malenia SGD
Regarding Malenia SGD’s termination condition (3), in Appendix E we demonstrate that this condition can be replaced with our simpler requirement of obtaining at least one gradient from every worker. With this modification, Malenia SGD remains optimal in the fixed-time regime (2) while becoming parameter-free, which eliminates the need for prior knowledge of $\sigma$ and $\varepsilon$. Under this parameter-free variant, the only difference between Malenia SGD and Ringleader ASGD lies in Phase 2: we perform updates asynchronously without discarding gradients, while Malenia SGD operates synchronously.
@@ 5.1 Iteration Complexity
The first lemma quantifies how much the objective function decreases at each iteration, accounting for both the standard gradient descent progress and the additional complexities introduced by asynchronous updates. {boxedlemma}[Descent Lemma; Proof in Section C.3] Under Assumptions 2.3 and 2.3, if the stepsize in Algorithm 1 satisfies $\gamma\leq\nicefrac{{1}}{{4L}}$, then the following inequality holds
@@
Next, similar to (Maranjyan et al., 2025d), we derive a lemma to bound the term involving the difference between current and old points. {boxedlemma}[Residual Estimation; Proof in Section C.4] Under Section 2.3, the iterates of Ringleader ASGD (Algorithm 1) with stepsize $\gamma\leq\nicefrac{{1}}{{4nL}}$ satisfy the following bound
@@ 5.2 Time Complexity
This time complexity result matches the lower bound derived by Tyurin & Richtárik (2024), thus making Ringleader ASGD the first asynchronous algorithm to achieve optimality.
@@ 7 Conclusion
Our analysis shows that Ringleader ASGD matches the optimal time complexity bounds established by Tyurin & Richtárik (2024). In contrast to the optimal but synchronous Malenia SGD method, Ringleader ASGD is asynchronous and requires no prior knowledge of problem parameters in the algorithm design, making it practical for real-world deployments.
Finally, with a minor modification, Ringleader ASGD also achieves optimality in the more general setting of arbitrarily varying computation times (Appendix B).
@@ Appendix A Related Work
Research on asynchronous stochastic gradient methods dates back to the seminal work of Tsitsiklis et al. (1986), and gained renewed momentum with the introduction of the HOGWILD! algorithm (Recht et al., 2011). HOGWILD! is fundamentally an asynchronous coordinate descent method: updates are performed lock-free with inconsistent reads and writes, and its convergence guarantees rely on sparsity assumptions that are rarely realistic in modern large-scale machine learning. Subsequent refinements of this paradigm include works by J Reddi et al. (2015); Zhao & Li (2016); Mania et al. (2017); Leblond et al. (2018); Nguyen et al. (2018); Zhou et al. (2018), but these works remain tied to the coordinate descent setting with inconsistent memory accesses, and thus differ substantially from our focus.
Closer to our setting are works where updates are based on gradients that are applied consistently. Early contributions, typically under the homogeneous-data assumption (all workers sample from the same distribution), include the work of Agarwal & Duchi (2011), who studied convex objectives, as well as later extensions to the non-convex case such as the work of Lian et al. (2015) and Dutta et al. (2018), the latter analyzing exponentially distributed computation times. Other relevant results in this line include those of Feyzmahdavian et al. (2016); Zheng et al. (2017); Arjevani et al. (2020); Feyzmahdavian & Johansson (2023), all of which assume fixed delays. More recently, delay-adaptive methods have been proposed, aiming to improve performance by down-weighting very stale gradients (Cohen et al., 2021; Koloskova et al., 2022; Mishchenko et al., 2022).
Particularly relevant to our work are asynchronous variants of SAGA. Leblond et al. (2018) developed a shared-parameter version in the spirit of HOGWILD!, while Glasgow & Wootters (2022) studied a distributed setting that employs full gradients, in contrast to our stochastic-gradient perspective.
A large body of recent work investigates asynchronous methods in federated learning (FL), where clients hold data from heterogeneous distributions. Notable contributions include works by Xie et al. (2019); Mishchenko et al. (2022); Koloskova et al. (2022); Wang et al. (2022a; b); Glasgow & Wootters (2022); Fraboni et al. (2023); Zhang et al. (2023); Wang et al. (2024); Islamov et al. (2024); Alahyane et al. (2025).
@@
There is another line of work that began with Tyurin & Richtárik (2024), who established lower bounds for parallel methods and proposed optimal synchronous algorithms together with an asynchronous counterpart. Several follow-up papers extended this semi-asynchronous framework to other settings (Tyurin et al., 2024a; b; Tyurin & Richtárik, 2024; Maranjyan et al., 2025a).
@@
Appendix B Arbitrarily Changing Computation Times

### v3

@@ 1 Introduction
A more promising approach is to shift from hardware solutions to algorithmic ones by adopting asynchronous optimization methods. These methods remove the need for synchronization, allowing fast workers to contribute updates without waiting for slower ones (Tsitsiklis et al., 1986; Recht et al., 2011; Agarwal and Duchi, 2011; Dean et al., 2012; Li et al., 2014). Despite their appeal, asynchronous methods are more difficult to analyze. In particular, a meaningful analysis would require studying time to convergence, rather than iteration complexity only. While iteration complexity is the traditional metric in optimization, it does not necessarily reflect real-world training speed in parallel settings: a method that performs more iterations may finish faster in wall-clock time if those iterations can be computed without waiting for slow workers. This distinction raises a fundamental question: among all parallel methods, which ones are provably fastest in theory? To make this question precise, we restrict our attention to smooth nonconvex problems and to stochastic first-order methods, encompassing algorithms with or without synchronization. This will be the only setting considered in this paper.
@@
Each column indicates whether a method satisfies the following desirable properties: Optimal: achieves the theoretical lower bound derived by Tyurin and Richtárik (2024b) for parallel first-order stochastic methods in heterogeneous data setting. No sync.: does not require synchronization and is therefore asynchronous. No idle workers: all workers remain busy without waiting, so computational resources are fully utilized. No discarded work: no computation is wasted, and no worker is stopped mid-computation. Our new method, Ringleader ASGD, is the first asynchronous method to achieve optimal time complexity, while also ensuring full resource utilization (no idle workers) and no discarded computations / work.
@@
Naive Minibatch SGD (Section 3.1) $\frac{L_{f}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{n}\frac{\sigma^{2}}{n\varepsilon}\right)$
IA2SGD (Wang et al., 2025) (Appendix E) $\frac{L_{\max}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{n}\frac{\sigma^{2}}{n\varepsilon}\right)$ (†)
Malenia SGD (Tyurin and Richtárik, 2024b) $\frac{L_{f}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{\mathrm{avg}}\frac{\sigma^{2}}{n\varepsilon}\right)$
Ringleader ASGD (new) (Algorithm 1; Section 5.2) $\frac{L\Delta}{\varepsilon}\left(\tau_{n}+\tau_{\mathrm{avg}}\frac{\sigma^{2}}{n\varepsilon}\right)$ (‡)
Lower Bound (Tyurin and Richtárik, 2024b) $\frac{L_{f}\Delta}{\varepsilon}\left(\tau_{n}+\tau_{\mathrm{avg}}\frac{\sigma^{2}}{n\varepsilon}\right)$
@@
Recently, Tyurin and Richtárik (2024b) studied this very regime, where they derived lower bounds. They then proposed two algorithms: Rennala SGD, designed for the homogeneous data setting, where all workers draw samples from the same distribution, and Malenia SGD, for the heterogeneous data setting, where data distributions differ across workers. They showed that both methods are optimal—achieving the lower bounds—and, perhaps surprisingly, both are synchronous (they periodically synchronize the workers). The key idea in both is to fully utilize the available computational resources by keeping workers continuously busy: each worker computes independently, and synchronization occurs only after a sufficient number of gradient computations have been accumulated.
At first, the result of Tyurin and Richtárik (2024b) suggested a rather pessimistic outlook for asynchronous methods: despite their practical appeal, they showed that existing asynchronous methods are not optimal and that the method achieving the lower bound is synchronous. This created the view that optimality is inherently tied to synchronization. However, this view was overturned by Maranjyan et al. (2025d), who, in the homogeneous data setting, introduced Ringmaster ASGD—the first asynchronous SGD method to achieve the same optimal time complexity as the synchronous Rennala SGD. Although both methods share the same theoretical guarantees, Ringmaster ASGD can be faster than Rennala SGD in practice, since it avoids synchronization and benefits from more frequent updates.
@@
To the best of our knowledge, the only method proven to be optimal in the heterogeneous data setting is the synchronous algorithm Malenia SGD of Tyurin and Richtárik (2024b), which, notably, does not rely on similarity assumptions. However, synchronization is a major bottleneck in practice: although synchronous and asynchronous methods can share the same theoretical complexity, asynchronous methods are often faster in practice because they avoid costly synchronization and benefit from more frequent updates, as demonstrated in the homogeneous case by Maranjyan et al. (2025d).
This raises a fundamental question: Is it possible to design an asynchronous method that requires no similarity assumptions while still achieving optimal time complexity? In this paper, we answer this question affirmatively by introducing Ringleader ASGD, the first asynchronous SGD method that achieves optimal time complexity [^1] Throughout the paper, we refer to our method as optimal. Formally, this holds whenever the constant LL—associated with our new smoothness-type assumption (Section 2.3)—is at most a constant factor larger than the smoothness constant $L_{f}$ used in the derived lower bounds (Tyurin and Richtárik, 2024b). See Table 1 for details. in the heterogeneous data setting. Importantly, Ringleader ASGD attains this without relying on restrictive similarity assumptions.
@@ 1.1 Contributions
  • Optimal asynchronous SGD under data heterogeneity. We prove that Ringleader ASGD (Algorithm 1) is, to the best of our knowledge, the first asynchronous method in the heterogeneous setting under the fixed computation model (2) matching the lower bounds for parallel methods of Tyurin and Richtárik (2024b), when the smoothness-type constant $L$ in Section 2.3 is within a constant factor of the smoothness $L_{f}$ used to obtain the lower bounds (Table 1). Importantly, Ringleader ASGD attains this without any similarity assumptions across clients’ data.
@@
  • Parameter-free design. In contrast to the optimal synchronous method Malenia SGD (Tyurin and Richtárik, 2024b), which requires prior knowledge of the gradient variance bound and target accuracy, our method operates in the fixed computation time model without such parameters (except for the stepsize, needed only to match the optimal theoretical rate). This makes it far more practical for real-world deployments, where these quantities are typically unknown or difficult to estimate. The same parameter-free improvement can also be extended to Malenia SGD, as we discuss in Appendix F.
  • Universal computation model. In Appendix C, we extend our analysis beyond the fixed computation time model to the general setting of arbitrarily varying computation times, accommodating virtually any computational behavior, including stochastic or adversarial patterns, while retaining optimality time complexity.
@@ 2.1 Worker Heterogeneity Model
We first focus on the case where workers have constant computation speeds, as this setting is more intuitive and serves as a foundational model for understanding the dynamics of asynchronous distributed optimization. The extension to arbitrary computation times is presented in Appendix C.
@@
We assume communication is infinitely fast (taking 0 seconds), both from workers to the server and from the server to workers [^3] Alternatively, one could define $\tau_{i}$ as the time required for a worker to both compute a gradient and communicate it to the server, while keeping server-to-worker communication infinitely fast. Our upper bounds would still hold under this formulation, but the lower bounds would no longer apply, so we use the simpler model.. This is a modeling choice—arguably the simplest one—and has been the standard assumption in prior work (Mishchenko et al., 2022; Koloskova et al., 2022; Tyurin and Richtárik, 2024b; Maranjyan et al., 2025d), even if not always stated explicitly. We further discuss the motivation and limitations of this abstraction in Appendix B. A related study by Tyurin et al. (2024b) considers the case where communication is non-negligible and proposes techniques to address it.
@@ 2.3 Assumptions
By convention, $L_{\phi}$ denotes the smallest such constant. Note that Section 2.3 is stronger than requiring $f$ itself to be $L_{f}$–smooth, yet weaker than all $f_{i}$ being $L_{f_{i}}$–smooth. The following lemma establishes the relation among the constants $_{f}, LL$, and $L_{f_{i}}$. {boxedlemma}[Smoothness Bounds; Proof in Section D.1] Suppose Section 2.3 holds with constant $L>0$. Then $f$ is $L_{f}$–smooth with $L_{f}\leq L$. Moreover, if each $f_{i}$ is $L_{f_{i}}$–smooth, then Section 2.3 is satisfied, and we have
@@ 3 Background and Motivation
In this section, we review relevant existing methods in distributed optimization and discuss their limitations to motivate our algorithmic design. We begin with Naive Minibatch SGD as the canonical synchronous baseline, then consider Malenia SGD (Tyurin and Richtárik, 2024b), the first synchronous SGD method to attain optimal time complexity. We then turn to the challenges of asynchronous approaches, focusing on the recent IA2SGD (Wang et al., 2025) and its limitations. Finally, we outline how these limitations can be addressed and introduce the core idea behind our algorithm.
@@ 3.2 Malenia SGD
Malenia SGD (Tyurin and Richtárik, 2024b) addresses the straggler problem of Naive Minibatch SGD by ensuring continuous worker utilization, rather than waiting for the slowest worker in each iteration. However, it is fundamentally a Minibatch SGD algorithm: in every iteration, it constructs an unbiased gradient estimator from multiple gradients and then performs a synchronous update. The key distinction lies in how the batch is collected—Malenia SGD gathers gradients asynchronously, allowing potentially multiple contributions from the same worker within a single iteration, unlike Naive Minibatch SGD.
@@
as shown by Tyurin and Richtárik (2024b). The key improvement over Naive Minibatch SGD is that the variance term $\sigma^{2}$ is now multiplied by $\tau_{\text{avg}}$ instead of $\tau_{n}$. Since $\tau_{\text{avg}}\ll\tau_{n}$ in computationally heterogeneous environments, Malenia SGD can potentially provide substantial speedup in highly heterogeneous regimes.
@@ 3.4 IA2SGD
The iteration complexity of this algorithm was established by Wang et al. (2025), and by a straightforward conversion to the fixed computation time model (see Appendix E), we obtain the corresponding time complexity
@@ 4.4 Comparison to Malenia SGD
Regarding Malenia SGD’s termination condition (3), in Appendix F we demonstrate that this condition can be replaced with our simpler requirement of obtaining at least one gradient from every worker. With this modification, Malenia SGD remains optimal in the fixed-time regime (2) while becoming parameter-free, which eliminates the need for prior knowledge of $\sigma$ and $\varepsilon$. Under this parameter-free variant, the only difference between Malenia SGD and Ringleader ASGD lies in Phase 2: we perform updates asynchronously without discarding gradients, while Malenia SGD operates synchronously.
@@ 5.1 Iteration Complexity
The first lemma quantifies how much the objective function decreases at each iteration, accounting for both the standard gradient descent progress and the additional complexities introduced by asynchronous updates. {boxedlemma}[Descent Lemma; Proof in Section D.3] Under Assumptions 2.3 and 2.3, if the stepsize in Algorithm 1 satisfies $\gamma\leq\nicefrac{{1}}{{4L}}$, then the following inequality holds
@@
Next, similar to the work by Maranjyan et al. (2025d), we derive a lemma to bound the term involving the difference between current and old points. {boxedlemma}[Residual Estimation; Proof in Section D.4] Under Section 2.3, the iterates of Ringleader ASGD (Algorithm 1) with stepsize $\gamma\leq\nicefrac{{1}}{{4nL}}$ satisfy the following bound
@@ 5.2 Time Complexity
Remark 5.1.
The optimality claim for Section 5.2 holds when the smoothness-type constant $L$ in Section 2.3 is within a constant factor of the smoothness $L_{f}$ used to derive the lower bound (Tyurin and Richtárik, 2024b) (Table 1).
Under this condition, the resulting time complexity matches the lower bound of Tyurin and Richtárik (2024b), making Ringleader ASGD the first asynchronous algorithm to achieve optimality under heterogeneous data.
@@ 7 Conclusion
Our analysis shows that Ringleader ASGD matches the optimal time complexity bounds established by Tyurin and Richtárik (2024b). In contrast to the optimal but synchronous Malenia SGD method, Ringleader ASGD is asynchronous and requires no prior knowledge of problem parameters in the algorithm design, making it practical for real-world deployments.
Finally, with a minor modification, Ringleader ASGD also achieves optimality in the more general setting of arbitrarily varying computation times (Appendix C).
@@ Appendix A Related Work
Research on asynchronous stochastic gradient methods dates back to the seminal work of Tsitsiklis et al. (1986), and gained renewed momentum with the introduction of the HOGWILD! algorithm (Recht et al., 2011). HOGWILD! is fundamentally an asynchronous coordinate descent method: updates are performed lock-free with inconsistent reads and writes, and its convergence guarantees rely on sparsity assumptions that are rarely realistic in modern large-scale machine learning. Subsequent refinements of this paradigm include works by J Reddi et al. (2015); Zhao and Li (2016); Mania et al. (2017); Leblond et al. (2018); Nguyen et al. (2018); Zhou et al. (2018), but these works remain tied to the coordinate descent setting with inconsistent memory accesses, and thus differ substantially from our focus.
Closer to our setting are works where updates are based on gradients that are applied consistently. Early contributions, typically under the homogeneous-data assumption (all workers sample from the same distribution), include the work of Agarwal and Duchi (2011), who studied convex objectives, as well as later extensions to the non-convex case such as the work of Lian et al. (2015) and Dutta et al. (2018), the latter analyzing exponentially distributed computation times. Other relevant results in this line include those of Feyzmahdavian et al. (2016); Zheng et al. (2017); Arjevani et al. (2020); Feyzmahdavian and Johansson (2023), all of which assume fixed delays. More recently, delay-adaptive methods have been proposed, aiming to improve performance by down-weighting very stale gradients (Cohen et al., 2021; Koloskova et al., 2022; Mishchenko et al., 2022).
Particularly relevant to our work are asynchronous variants of SAGA. Leblond et al. (2018) developed a shared-parameter version in the spirit of HOGWILD!, while Glasgow and Wootters (2022) studied a distributed setting that employs full gradients, in contrast to our stochastic-gradient perspective.
A large body of recent work investigates asynchronous methods in federated learning (FL), where clients hold data from heterogeneous distributions. Notable contributions include works by Xie et al. (2019); Mishchenko et al. (2022); Koloskova et al. (2022); Wang et al. (2022a; b); Glasgow and Wootters (2022); Fraboni et al. (2023); Zhang et al. (2023); Wang et al. (2024); Islamov et al. (2024); Alahyane et al. (2025).
@@
There is another line of work that began with Tyurin and Richtárik (2024b), who established lower bounds for parallel methods and proposed optimal synchronous algorithms together with an asynchronous counterpart. Several follow-up papers extended this semi-asynchronous framework to other settings (Tyurin et al., 2024a; b; Tyurin and Richtárik, 2024a; Maranjyan et al., 2025a).
@@
Appendix B The Computation-Only Model and the Role of Asynchrony
This section clarifies the scope of our modeling assumptions and the specific phenomenon our results target. Asynchrony is designed to eliminate waiting time: the idle time that arises when workers have heterogeneous computation speeds or experience straggling behavior due to hardware stalls, load imbalance, or even network delays. A key point is that the goal of asynchrony is not to reduce communication cost, but to ensure that slow or delayed workers do not force faster workers to remain idle.
Critically, even in the simplest setting—homogeneous data and no communication cost—it was unknown until very recently whether an asynchronous SGD method could match the optimal synchronous rate. In fact, existing asynchronous methods were shown to be worse than the optimal synchronous algorithm in this basic regime (Tyurin and Richtárik, 2024b). The recent work of Maranjyan et al. (2025d) resolved this foundational case for the first time by showing that, under homogeneous data, an asynchronous method can achieve the same optimal time complexity as the synchronous method Rennala SGD (Tyurin and Richtárik, 2024b). Our work extends this understanding to the significantly more challenging heterogeneous-data setting and shows that asynchrony can solve the problem it is designed for—and that it can do so optimally.
Asynchrony and stragglers.
Stragglers may be caused by slow computation, device variability, or even communication delays. From the server’s perspective, a worker whose gradient arrives late because it is still communicating appears identical to one that is slow at computing. Asynchrony ensures that such delays—regardless of source—do not block progress: fast workers continue contributing updates while slow or delayed workers catch up. This ability to eliminate idle time is precisely the purpose of asynchronous methods.
Asynchrony does not reduce communication cost.
Although asynchrony prevents waiting during communication-induced delays, it does not reduce the communication overhead itself. Reducing communication cost requires orthogonal techniques such as gradient compression, sparsification, quantization, or local-update schemes (e.g., FedAvg (McMahan et al., 2016)). For example, Tyurin et al. (2024b) explicitly study communication-aware training and use compression-based methods to reduce communication time—demonstrating that communication efficiency is a separate algorithmic axis that must be combined with, rather than replaced by, asynchrony. Thus, asynchrony resolves the waiting problem caused by delays, but not the communication-cost problem; addressing the latter requires additional mechanisms.
Why communication is not modeled explicitly here.
For these reasons, we adopt the standard computation-only model used in essentially all theoretical works on asynchronous SGD (Mishchenko et al., 2022; Koloskova et al., 2022; Tyurin and Richtárik, 2024b; a; Maranjyan et al., 2025d), which is also the model under which the lower bounds of Tyurin and Richtárik (2024b) are derived. Our claims of optimal time complexity therefore refer to this shared and well-established model. Studying asynchrony under explicit communication cost—where it must interact with compression, local updates, or buffering—requires new lower bounds and a different theoretical framework, and is beyond the scope of this work.
Appendix C Arbitrarily Changing Computation Times